OpenStudy (anonymous):

Conjecture a formula for un=un-1-2 where u0=4

OpenStudy (anonymous):

4-2*n

OpenStudy (anonymous):

are yu sure that's typed right?

OpenStudy (anonymous):

\[u _{n}=u _{n-1}+d\]

OpenStudy (anonymous):

oh ok. so,\[u_{0} = 4 ?\]

OpenStudy (anonymous):

\[u _{_{0}}=4\]

OpenStudy (anonymous):

than it is 4+n*d

OpenStudy (anonymous):

is this that you need?

OpenStudy (anonymous):

I saw that the equation would be y=-2x +4 but need a general solution in terms of recurrence

OpenStudy (anonymous):

un=un−1+d if this is the general formula that the answer should be u0 + n times d

OpenStudy (anonymous):

y=-2x +4 this is the same as I wrote first just written as a function

OpenStudy (anonymous):

yes but I need it in terms of previous

OpenStudy (anonymous):

Use iteration (by hand or with your calculator) to explore the first few terms of each sequence. ii. On the basis of (i), conjecture a formula for a solution u(n) of the general affine system with r=1: (I) u0=c (R) un=un-1+d for n > 0